# Pseudodifferential damping estimates and stability of relaxation shocks

### v3

@@ 1.1. Main results
Along the way, we discuss in Sections 2.5.1, 3.4.1, 4.1, 6.1, and 7.3 the changes needed for extension to the discontinuous case, in particular sketching the treatment in 1- and multi-D of arbitrary-amplitude shock profiles with a single subshock discontinuity, of the types arising for discontinuous hydraulic shock waves of the Saint Venant equations [54, 11, 55]. See Proposition 3.9, Corollary 4.3, and the treatment in Theorem 6.4 of a multi-d turning point case scenario noted by Erpenbeck [8, 27, 61] in the case of detonations that as shown in [55] is relevant for discontinuous hydraulic shock profiles of the Saint Venant equations of inclined shallow-water flow. See Theorem 6.4.
@@ 2.2. Paralinearization
To remove extraneous symmetry assumptions and connect directly to the linearized resolvent equation, we perform a pseudodifferential analysis, replacing $\partial_{x_{j}}$, $j=2,\dots,d$ and $\partial_{t}$ by their symbols $i\eta_{j}$ and $i\tau$. See [42, Appendix B] for a description of the relevant tools, phrased in the paradifferential calculus of Bony; different from the setting of [42], there is no additional parameter here such as viscosity or relaxation constant, simplifying slightly the situation. This has the added advantage, in the case of planar traveling-wave profiles containing discontinuous subshocks, of putting things in a common framework with the finite-time stability analysis of discontinuous shock problems pioneered in [24, 32, 33, 38], where frequency-dependent (in time) estimates are seen to be necessary for multi-D stability estimates.
@@
where $\Gamma=\Gamma(w,\eta,\lambda)$ is a frequency-dependent boundary condition obtained from the linearized Rankine-Hugoniot jump conditions about $w$, with estimates for the subshock front perturbation $\psi$ recoverable from $v$. For further details, see, e.g., [38, 54], or [2, Appendix].
@@ 3.1. Symmetrizers
We begin by repeating essentially verbatim the generalization of Kreiss’ method to variable-coefficient problems in [41, §2.1], there stated for the discontinuous case. In the simpler smooth case, the spatial domain is extended to $(-\infty,\infty)$ and (3.4) dropped. Recall now the essence of the “method of symmetrizers” as it applies to general boundary value problems
@@ 3.2.1. The quantitative estimates of Lannes
For further details, see the commutator estimates of [29, Thm. 7, p. 529], the quantity $H^{t_{0}}_{ref}$ defined in [29, Eq. (4.7), p. 519], and the quantities $n$ and $N$ on which $H^{t_{0}}_{ref}$ depends, defined in [29, Eqs. (2.2) and (2.4), pp. 502 and 503]; see also [29, Thms. 1 and 2, pp. 514 and 515] and [29, Cor. 30, p. 515] bounding the action of operators in Sobolev norms in terms of Sobolev norms of their symbols. For the main, $H^{s}$ (rather than $H^{s-1}$ of gradient) bounds, see also the alternative, more streamlined treatment of Sroczinski in [51, §2], focused entirely on Sobolev-based bounds. For general discussions of pseudo- and paradifferential calculus, see [3, Appendix C] and [40].
@@ 3.3.1. The symmetric case
This may be compared with the standard ($G$ rather than $\tilde{G}$) symmetrizer construction of [43, Appendix E] for hyperbolic-parabolic, or “viscous” conservation laws
@@ 3.4.1. The discontinuous case
with $C>0$ sufficiently large, we obtain (3.4), exactly as in the treatment by Kreiss of the constant coefficient case with spectral gap. (See first part of proof of Main Theorem 4 [24, top, p. 292].)
@@ 4. The large-amplitude 1-D noncharacteristic hyperbolic case
we may thus appeal to the tracking/reduction lemma of [37, §3.2] to conclude the existence of an ${\rm Id}+O(\varepsilon)$ coordinate change converting (4.11) to exact diagonal form
@@ 5.1. Kreiss’ glancing
for $\tilde{\gamma}$ bounded and $\varepsilon$ sufficiently small. For the general case accounting for neglected terms, see, e.g., [24, §4], [38], [42, Appendix A], or [19, §5].
@@ 5.2. Variable coefficients and Airy-type turning points
For $a^{\prime}(0)=1$, the point $x=0$ in (5.6), corresponding to the special point $x_{*}$ in the original equations, under the assumptions (5.7) can be recognized as an “Airy-type” turning point; cf. [53, 27]. That is, glancing points in $\lambda$ are replaced for variable coefficients by turning points in $x$. Note in (5.6) that $\Im\lambda$ has been absorbed in shifts of $a$, $x$, hence $\tilde{\tau}$ in (5.2) effectively replaced by $a(x)/\varepsilon$. An important difference from the constant-coefficient case is that $ia(x)/\varepsilon$ is more singular than the former term $\tau$, requiring additional care in the symmetrizer construction. In particular, the classical construction of Kreiss can be expected a priori to succeed only in a small neighborhood of $x=0$, where $|a(x)|/\varepsilon$ is small compared to $1/\varepsilon$. (The setting of Kreiss, translated to our coordinates, is $\tau=O(1)$; what is needed for his arguments is $\tau\ll 1/\varepsilon$ [24, §4].)
@@ 6. A basic large-amplitude damping result in multi-D
1. The reduction to traceless form (6.6) may be regarded as a variant of [24, Lemma 4.3] in the special case of a Jordan block of dimension $s=2$, designed to highlight the connection with Airy’s equation and the analyses of [8, 27, 55]. As can be seen by the details of the argument, form (5.6) is not needed on the whole line, but only in a vicinity of the turning point $x_{*}$, allowing treatment of more general cases with further care.
2. The crucial construction of the family of symmetrizers $S_{\alpha}$ likewise reflects the key observation of [24, Lemma 4.4] for Jordan blocks of arbitrary dimension $s\geq 2$ that there is a degree of freedom in the choice of symmetrizers allowing arbitrarily large constants in coordinates $2,\dots,s$, allowing absorbtion of remaining error terms.
3. The block structure condition for Airy point at infinity, though phrased here as an additional assumption, is in the hyperbolic constant multiplicity case a consequence of high-frequency dissipativity of the end states, which is implied by our standard assumption of high-frequency resolvent bound (1.3). See [24, Lemma 2.7, Eqn. (2.21)] in the stictly hyperbolic case, [39] in the general constant-multiplicity case.
@@ 6.1.4. Coinciding subshock and turning point
generalizing (5.14). For, as noted in [24, Lemma 4.1], there is a unique one-parameter family of symmetrizers $D$ of the principal Jordan block part
@@ 7.1. Short-time theory
(7.3) $\|v(t)\|_{H^{s}}^{2}\leq C\|v(0)\|_{H^{s}}^{2}+C\int_{0}^{t}\|f(\tau)\|_{H^{s}}^{2}\,d\tau$
for all $t\geq 0$, a result holding also in the original coordinates.
@@
with $\|v(t)\|_{H^{s}}$ continuous. Moreover, the solution continues so long as the Lipshitz norm of $v$ remains bounded and the $L^{\infty}$ norm of $v$ remains sufficiently small.
7.2. Truncation to finite time
Proposition 7.1 seems interesting as a frequency-dependent version of Friedrichs symmetric hyperbolic theory. Equally important for our purposes is the underlying estimate (7.3), and the resulting bounds for solutions defined on $[0,T]$ of
@@ 7.2. Truncation to finite time
Assuming (A2)-(A3), for a solution $w=\bar{w}+v$ of (2.4) on $[0,T]$, satisfying (2.9) and with $v$ sufficiently small in $H^{s}([0,T]\times\mathbb{R}^{d})$, there holds the nonlinear damping estimate (2.3).
@@
Setting $\tilde{v}:=\chi_{1}(t)\chi_{T}(t)v$ and $\tilde{f}=\partial_{x_{1}}\tilde{v}-G\tilde{v}$, where $\chi_{1}$ is a $C^{\infty}$ cutoff function equal to one for $t\geq\tau$ and zero for $t\leq 0$ and $\chi_{T}$ is a $C^{\infty}$ cutoff function equal to one for $t\leq T-\tau$ and zero for $t\geq T$, and applying (2.11), we obtain
@@
Adding to this equation the equations (7.4) and (7.5), bounding the term $C\int_{T-2\tau}^{T-\tau}\|v(t)\|_{H^{s}}^{2}dt$ on the righthand side of (7.5) using again (7.6), and rearranging, we obtain the result. ∎
@@
Conclusion: Theorem 7.2 completes our goal for smooth waves, recovering the central estimate in the classical damping case, sufficient to close a nonlinear iteration and establish stability, provided that there hold a pseudodifferential damping estimate (2.9).

### v4

@@ 1.1. Main results
Along the way, we discuss in Sections 2.5.1, 3.4.1, 4.1, 6.1, and 7.4 the changes needed for extension to the discontinuous case, in particular sketching the treatment in 1- and multi-D of arbitrary-amplitude shock profiles with a single subshock discontinuity, of the types arising for discontinuous hydraulic shock waves of the Saint Venant equations [59, 12, 60]. See Proposition 3.9, Corollary 4.3, and the treatment in Theorem 6.4 of a multi-d turning point case scenario noted by Erpenbeck [9, 30, 67] in the case of detonations that as shown in [60] is relevant for discontinuous hydraulic shock profiles of the Saint Venant equations of inclined shallow-water flow.
@@ 2.2. Paralinearization
To remove extraneous symmetry assumptions and connect directly to the linearized resolvent equation, we perform a pseudodifferential analysis, replacing $\partial_{x_{j}}$, $j=2,\dots,d$ and $\partial_{t}$ by their symbols $i\eta_{j}$ and $i\tau$. See [45, Appendix B] for a description of the relevant tools, phrased in the paradifferential calculus of Bony; different from the setting of [45], there is no additional parameter here such as viscosity or relaxation constant, simplifying slightly the situation. This has the added advantage, in the case of planar traveling-wave profiles containing discontinuous subshocks, of putting things in a common framework with the finite-time stability analysis of discontinuous shock problems pioneered in [27, 35, 36, 41], where frequency-dependent (in time) estimates are seen to be necessary for multi-D stability estimates.
@@
where $\Gamma=\Gamma(w,\eta,\lambda)$ is a frequency-dependent boundary condition obtained from the linearized Rankine-Hugoniot jump conditions about $w$, with estimates for the subshock front perturbation $\psi$ recoverable from $v$. For further details, see, e.g., [41, 59], or [3, Appendix].
@@ 3.1. Symmetrizers
We begin by repeating essentially verbatim the generalization of Kreiss’ method to variable-coefficient problems in [44, §2.1], there stated for the discontinuous case. In the simpler smooth case, the spatial domain is extended to $(-\infty,\infty)$ and (3.4) dropped. Recall now the essence of the “method of symmetrizers” as it applies to general boundary value problems
@@ 3.2.1. The quantitative estimates of Lannes
For further details, see the commutator estimates of [32, Thm. 7, p. 529], the quantity $H^{t_{0}}_{ref}$ defined in [32, Eq. (4.7), p. 519], and the quantities $n$ and $N$ on which $H^{t_{0}}_{ref}$ depends, defined in [32, Eqs. (2.2) and (2.4), pp. 502 and 503]; see also [32, Thms. 1 and 2, pp. 514 and 515] and [32, Cor. 30, p. 515] bounding the action of operators in Sobolev norms in terms of Sobolev norms of their symbols. For the main, $H^{s}$ (rather than $H^{s-1}$ of gradient) bounds, see also the alternative, more streamlined treatment of Sroczinski in [55, §2], focused entirely on Sobolev-based bounds. For general discussions of pseudo- and paradifferential calculus, see [4, Appendix C] and [43].
@@ 3.3.1. The symmetric case
This may be compared with the standard ($G$ rather than $\tilde{G}$) symmetrizer construction of [46, Appendix E] for hyperbolic-parabolic, or “viscous” conservation laws
@@ 3.4.1. The discontinuous case
with $C>0$ sufficiently large, we obtain (3.4), exactly as in the treatment by Kreiss of the constant coefficient case with spectral gap. (See first part of proof of Main Theorem 4 [27, top, p. 292].)
@@ 4. The large-amplitude 1-D noncharacteristic hyperbolic case
we may thus appeal to the tracking/reduction lemma of [40, §3.2] to conclude the existence of an ${\rm Id}+O(\varepsilon)$ coordinate change converting (4.11) to exact diagonal form
@@ 5.1. Kreiss’ glancing
for $\tilde{\gamma}$ bounded and $\varepsilon$ sufficiently small. For the general case accounting for neglected terms, see, e.g., [27, §4], [41], [45, Appendix A], or [22, §5].
@@ 5.2. Variable coefficients and Airy-type turning points
For $a^{\prime}(0)=1$, the point $x=0$ in (5.6), corresponding to the special point $x_{*}$ in the original equations, under the assumptions (5.7) can be recognized as an “Airy-type” turning point; cf. [58, 30]. That is, glancing points in $\lambda$ are replaced for variable coefficients by turning points in $x$. Note in (5.6) that $\Im\lambda$ has been absorbed in shifts of $a$, $x$, hence $\tilde{\tau}$ in (5.2) effectively replaced by $a(x)/\varepsilon$. An important difference from the constant-coefficient case is that $ia(x)/\varepsilon$ is more singular than the former term $\tau$, requiring additional care in the symmetrizer construction. In particular, the classical construction of Kreiss can be expected a priori to succeed only in a small neighborhood of $x=0$, where $|a(x)|/\varepsilon$ is small compared to $1/\varepsilon$. (The setting of Kreiss, translated to our coordinates, is $\tau=O(1)$; what is needed for his arguments is $\tau\ll 1/\varepsilon$ [27, §4].)
@@ 6. A basic large-amplitude damping result in multi-D
1. The reduction to traceless form (6.6) may be regarded as a variant of [27, Lemma 4.3] in the special case of a Jordan block of dimension $s=2$, designed to highlight the connection with Airy’s equation and the analyses of [9, 30, 60]. As can be seen by the details of the argument, form (5.6) is not needed on the whole line, but only in a vicinity of the turning point $x_{*}$, allowing treatment of more general cases with further care.
2. The crucial construction of the family of symmetrizers $S_{\alpha}$ likewise reflects the key observation of [27, Lemma 4.4] for Jordan blocks of arbitrary dimension $s\geq 2$ that there is a degree of freedom in the choice of symmetrizers allowing arbitrarily large constants in coordinates $2,\dots,s$, allowing absorbtion of remaining error terms.
3. The block structure condition for Airy point at infinity, though phrased here as an additional assumption, is in the hyperbolic constant multiplicity case a consequence of high-frequency dissipativity of the end states, which is implied by our standard assumption of high-frequency resolvent bound (1.3). See [27, Lemma 2.7, Eqn. (2.21)] in the stictly hyperbolic case, [42] in the general constant-multiplicity case.
@@ 6.1.4. Coinciding subshock and turning point
generalizing (5.14). For, as noted in [27, Lemma 4.1], there is a unique one-parameter family of symmetrizers $D$ of the principal Jordan block part
@@ 7.1. Short-time theory
(7.3) $\|v(t)\|_{H^{s}}^{2}\leq e^{Ct}\|v(0)\|_{H^{s}}^{2}+C\int_{0}^{t}e^{C(t-\tau)}\|f(\tau)\|_{H^{s}}^{2}\,d\tau$
for all $t\geq 0$, so long as the Lipshitz norm of $v$ remains bounded, a result holding also in the original coordinates.
@@
with $\|v(t)\|_{H^{s}}$ upper semicontinuous. Moreover, the solution continues so long as the Lipshitz norm of $v$ remains bounded and the $L^{\infty}$ norm of $v$ remains sufficiently small.
7.2. Linear bounds
Proposition 7.1 seems interesting as a frequency-dependent version of Friedrichs symmetric hyperbolic theory. In the linear setting, it yields global existence with bounded exponential growth, furnishing a starting point from which we may verify existence of a Laplace transform, in order to obtain the following sharpened bounds. Consider the linear homogeneous version
(7.4) $(\partial_{t}-L)v=0$
of (7.1).
Proposition 7.2.
Assume along with (A2)-(A3) that the linearized operator $L$ have no spectra with positive real part. Then, assuming the high-frequency resolvent estimate (1.3) (as follows from a linear damping estimate), the solution of (7.4) satisfies for any $\theta>0$, and some $C=C(\theta>0$,
(7.5) $\int_{0}^{+\infty}e^{-2\theta s}\|v(t)\|_{H^{s}}^{2}ds\leq C\|v(0)\|_{H^{s}}^{2}.$
Proof.
Consider $\tilde{v}:=\chi(t)v$, where $\chi$ is a $C^{\infty}$ cutoff function vanishing at $t=0$ and $1$ for $t\geq 1$. Note that $\tilde{v}$ satisfies $(\partial_{t}-L)\tilde{v}=\chi^{\prime}(t)\tilde{v}$, with $\chi^{\prime}$ supported on $0\leq t\leq 1$, and $\|v(t)\|_{H^{s}}\leq C\|v(0)\|_{H^{s}}$ for $0\leq t\leq 1$. By Parseval’s inequality, we thus obtain (7.5) for all $\theta$ for which the inverse Laplace transform formula holds in $L^{2}$ sense in time, in particular for $\Re\lambda\geq C$, with $C>0$ sufficiently large, as in (7.3). Using the resolvent identity as in standard semigroup theory, together with the assumed uniform high-frequency estimate (1.3), and the assumed boundedness of the resolvent on compact frequency sets with $\Re\lambda>0$, we may move the contour to $\Re\lambda=\theta$ for any $\theta>0$, so long as $\|Lv(0)\|\leq C$, yielding the result for $v(0)\in D(L)$. The result for general $v(0)$ then follows by a density argument, using boundedness/continuity with respect to initial data of the linearized solution operator. For smooth solutions, one may alternatively apply Prüss theorem [52] to obtain the same result. ∎
Equally important for our purposes is the underlying estimate (7.3), and the resulting bounds for solutions defined on $[0,T]$ and any finite $\tau>0$, of
@@ 7.2. Linear bounds
7.3. Truncation to finite time
We are now ready to establish our main, finite-time result.
@@ 7.3. Truncation to finite time
Assuming (A2)-(A3), for a solution $w=\bar{w}+v$ of (2.4) on $[0,T]$, for which (i) there exist symmetrizers yielding (2.9), (ii) the linearized operator $L$ about $\bar{w}$ has no spectra with positive real part, and (iii) $v$ is sufficiently small in $H^{s}([0,T]\times\mathbb{R}^{d})$, there holds for some $\eta>0$ sufficiently small the nonlinear damping estimate (2.3).
@@
First, notice that we may easily reduce to the case that $f$ is supported on $t\leq T$, by multiplying $f$ by a $C^{\infty}$ cutoff $\chi(t)$ equal to $1$ for $t=T-\tau$ and 0 for $t=T$. Denoting the resulting modified solution as $\tilde{v}$ and applying (2.11), we obtain
@@
Adding to this equation the equation (7.7), bounding the term $C\int_{T-2\tau}^{T-\tau}\|v(t)\|_{H^{s}}^{2}dt$ on the righthand side of (7.7) using again (7.8), and rearranging, we obtain the result, assuming that it holds for $C^{\infty}$ $f$ supported on $t\leq T$.
Next, consider $C^{\infty}$ $f$ supported on $0\leq t\leq T$, and $v$ vanishing at $t=0$, and introduce the modified function $v^{m}$ defined as follows. For $0\leq t\leq T$, take $v^{m}=v$. For $t\geq T$, let $v^{m}$ satisfy the homogeneous linear equation about profile $\bar{w}_{x}$. For $T\leq t\leq T+1$, let $v^{m}$ satisfy a homogeneous equation obtained by $C^{\infty}$ interpolation between $\bar{w}$ and $\bar{w}+v$ in the arguments of the coefficients of (2.4). Finally, set
(7.9) $v^{\dagger}:=(1+\chi(t)(e^{-\tilde{\theta}(t-T)}-1)v^{m},$
where $\chi$ is a $C^{\infty}$ cutoff function vanishing for $0\leq t\leq T$ and identically one for $t\geq T+1$. Then, we find that $v^{\dagger}$ satisfies (2.4), with modified coeffients as described above, and with modified forcing $f^{\dagger}$ defined as $f$ for $0\leq t\leq T$, $\tilde{\theta}v^{\dagger}$ for $t\geq T+1$, and bounded for $T\leq t\leq T+1$ by
(7.10) $\xi^{\prime}(e^{-\tilde{\theta}(t-T)}-1)+\tilde{\theta}\chi(e^{-\tilde{\theta}(t-T}-1)=O(\tilde{\theta})\|v^{m}(t)\|_{H^{s}}=O(\tilde{\theta}\|v(T)\|_{H^{s}},$
where the final inequality follows for $\|v(T)\|_{H^{s}}$ sufficiently small by short-time existence theory applied on the time interval $T\leq t\leq T+1$.
By Proposition 3.3, there exist symmetrizers yielding (2.9) for this interpolant as well. Moreover, by linear bounds (7.5), the Laplace transform of $v^{\dagger}$ is well-defined for $\Re\lambda\geq\gamma=-\eta$ for $\eta>0$ sufficiently small relative to $\tilde{\theta}$. Applying now (2.10), we thus obtain
(7.11) $\int_{0}^{+\infty}e^{2\gamma(T-t}\|v^{\dagger}(t)\|_{H^{s}}^{2}dt\leq C_{2}\int_{0}^{+\infty}e^{2\gamma(T-t)}\big(\|f^{\dagger}(t)\|_{H^{s}}^{2}+\|v^{\dagger}(t)\|_{L^{2}}\big)dt.$
By (7.10), using again (7.7), the contribution of $f^{\dagger}$ from $T\leq t\leq T+1$ may be absorbed in the lefthand side for $\tilde{\theta}$ sufficiently small. For $t\geq T+1$ on the other hand, $\|f^{\dagger}\|_{H^{s}}=\tilde{\theta}\|v^{\dagger}\|_{H^{s}},$ and so can again be absorbed on the lefthand side of (7.11) for $\tilde{\theta}$ sufficiently small.
Finally, by linear estimate (7.5), the contribution of $\|v^{\dagger}(t)\|^{2}$ for $t\geq T+1$ may be bounded by $C\|\dagger v(T+1)\|_{L^{2}}^{2}$, while by short-time existence theory the contribution from $T\leq t\leq T+1$ may be bounded by $C\|v(T)\|_{L^{2}}^{2}$, each of which by (7.7) may be bounded by $C\int_{T-1}^{T}\|v(t)\|_{L^{2}}^{2}dt$, hence absorbed in the $L^{2}$ contribution for $0\leq t\leq T$ on the righthand side of (7.11).
Removing all of these absorbable terms reduces the righthand side of (7.11) to an integral from 0 to $T$. Discarding the terms for $t\geq T$ on the lefthand side and multiplying by $e^{2\gamma T}$ thus gives
$\int_{0}^{T}e^{2\gamma(T-t}\|v^{\dagger}(t)\|_{H^{s}}^{2}dt\leq C_{2}\int_{0}^{+\infty}e^{2\gamma(T-t)}\big(\|f^{\dagger}(t)\|_{H^{s}}^{2}+\|v^{\dagger}(t)\|_{L^{2}}\big)dt,$
establishing (2.11) for the key case of $C^{\infty}$ $f$ supported on $t\geq 0$ and $v$ vanishing at initial time $t=0$. From (2.11), we readily obtain (2.3) from the bound
$\|v(T)\|_{H^{s}}^{2}\leq C\int_{T-1}^{T}\|v(t)\|_{H^{s}}^{2}\,dt$
similar to (7.7), established by short time bound $\|v(T)\|_{H^{s}}\leq C\int_{T-1}^{T}\|v(t)\|_{H^{s}}^{2}\,dt$ together with Jensen’s inequality.
The case of nonvanishing initial data or forcing is now straightforward, introducing the modified function $v\sharp:=\chi(t)v$ with $\chi$ a smooth cutoff function that is identically zero for $t\leq 0$ and identically one for $t\geq 1$, and using (7.6). This satisfies (2.4) with $C^{\infty}$ forcing $f^{\sharp}:=\chi f+\chi^{\prime}v$ vanishing for $t\leq 0$ and $v^{\sharp}$ vanishing at $t=0$, agreeing with $f$ and $v$ for $t\geq 1$. Estimates for $0\leq t\leq 1$ may be obtained from short-time theory, completing the proof. ∎
@@
Conclusion: Theorem 7.3 completes our goal for smooth waves, recovering the central estimate in the classical damping case, sufficient to close a nonlinear iteration and establish stability, provided there exist symmetrizers yielding a pseudodifferential damping estimate (2.9).
Remark 7.4.
Alternatively, the theorem may be phrased without reference to symmetrizers by assuming there exist damping estimates (2.9) for both the paralinearized problem and the modified problem obtained as described above by interpolating coefficient arguments toward the linearized problem over the time interval $T\leq t\leq T+1$.
Remark 7.5.
As for the symmetrizer estimates earlier on, the step of truncation requires additional care in the damping case compared to classical arguments for short-time well-posedness, in which latter case error terms need only be shown to be order one and not small in order to be absorbed. In particular, we seem to require the additional assumption (ii) that the linearized operator $L$ about the (exact) traveling wave have no spectra of positive real part. It is an interesting open question whether this is a technical artifact or indeed necessary for the result. Examination of the linear setting suggests that presence of essential spectra with positive real part may indeed present a problem, an issue connected with causality in the inverse Laplace transform formula. However, in any case condition (ii) is no restriction for the use in nonlinear stability arguments for which the damping estimates are intended, since absence of positive real part spectra of $L$ is a necessary condition for the linearized stability estimates on which such arguments are based.